OpenStudy (anonymous):

What is the equation of the line which includes points (2, 3) and (0, 10)? A y=-7/2x-10 B y=-2/7x+10 C y=-7/2x+10 D y=2/7x+10

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Standard: y = mx+b m is the slope b is the y-intercept (the value of y when x=0)

OpenStudy (anonymous):

where m=(y2-y1)\(x2-x1)=(10-3)\(2-0)=7\2 ,therefore :y=7\2x+10 ,but since there is no choise =that,so we will take a negative sign as a common so it will be -(7\2x-10) which is (A)

OpenStudy (anonymous):

if u found this answer is suitable and a good one ,plz choose as a good answer ,ty "_)